### v1

# SwiftTry: Fast and Consistent Video Virtual Try-On with Diffusion Models

Abstract
Given an input video of a person and a new garment, the objective of this paper is to synthesize a new video where the person is wearing the specified garment while maintaining spatiotemporal consistency. While significant advances have been made in image-based virtual try-ons, extending these successes to video often results in frame-to-frame inconsistencies. Some approaches have attempted to address this by increasing the overlap of frames across multiple video chunks, but this comes at a steep computational cost due to the repeated processing of the same frames, especially for long video sequence. To address these challenges, we reconceptualize video virtual try-on as a conditional video inpainting task, with garments serving as input conditions. Specifically, our approach enhances image diffusion models by incorporating temporal attention layers to improve temporal coherence. To reduce computational overhead, we introduce ShiftCaching, a novel technique that maintains temporal consistency while minimizing redundant computations. Furthermore, we introduce the TikTokDress dataset, a new video try-on dataset featuring more complex backgrounds, challenging movements, and higher resolution compared to existing public datasets. Extensive experiments show that our approach outperforms current baselines, particularly in terms of video consistency and inference speed. Data and code are available at https://github.com/VinAIResearch/swift-try.
Introduction
Figure 1: Results of our SwiftTry compared with those of ViViD (Fang et al. 2024), the previous method for video try-on. Our method maintains detail and consistency in the texture of the garments and runs over 60% faster.
Video virtual try-on is an emerging research area (Chen et al. 2021; Rogge et al. 2014; Pumarola et al. 2019; Dong et al. 2019b; Kuppa et al. 2021; Zhong et al. 2021; Jiang et al. 2022; He et al. 2024; Xu et al. 2024b; Fang et al. 2024; Zheng et al. 2024) with significant potential in fashion and e-commerce. The ability to realistically visualize how a garment looks on a person in a video could transform online shopping. However, despite recent progress in image-based virtual try-on (He, Song, and Xiang 2022; Choi et al. 2021; Lee et al. 2022; Xie et al. 2023; Zhu et al. 2023; Kim et al. 2023), extending these capabilities to video remains challenging due to the need for spatiotemporal consistency and the high computational costs of processing long sequences.
A significant challenge in video virtual try-on is balancing the need for temporal coherence with the computational demands of processing long video sequences. Previous methods (Xu et al. 2024b; He et al. 2024; Fang et al. 2024) often struggle with temporal inconsistencies, leading to visual artifacts and flickering between frames, which undermines the realism of the virtual try-on experience. Additionally, the high computational cost of rendering high-quality results over extended sequences hampers the practicality of these approaches for real-world applications.
Another challenge is lacking an evaluation dataset. The first public video try-on dataset VVT (Dong et al. 2019b) only covers basic pattern garments, form-fitting T-shirts, uniform backgrounds, static camera angles, and repetitive human motions. Recently, ViViD (Fang et al. 2024) released the first practical dataset for video virtual try-on; however, it struggles to handle in-the-wild scenarios, such as complex movements and diverse backgrounds, making it challenging to meet the demands of practical applications. Moreover, one reason for the poor quality video try-on result is primarily due to masks extracted using human parsing segmentation (Li et al. 2020) applied on each frame of the video.
In this paper, we tackle these challenges by introducing two key contributions. First, we present a new high-quality dataset named TikTokDress consisting of 817 videos which is specifically designed for training and evaluating video virtual try-on models. This dataset features realistic scenes, diverse garment types, and complex movements, providing a robust foundation for advancing research in this field. Second, we introduce a novel video virtual try-on framework named SwiftTry, as shown in Fig. 1 which significantly reduces the computational cost of processing long video sequences while maintaining temporal consistency. Our framework is inspired by state-of-the-art diffusion-based image virtual try-on methods (Kim et al. 2023; Xu et al. 2024a; Choi et al. 2024) and incorporates temporal attention in the UNet architecture to train on video try-on data. During inference, we introduce a new technique called ShiftCaching, which ensures temporal coherence and smooth transitions between video clips while also reducing redundant computation compared to previous methods. Extensive experimental results indicate that our proposed SwiftTry framework, incorporating the aforementioned techniques, substantially surpasses other video virtual try-on methods in performance.
In summary, the contributions of our work are as follows:
  • We propose a new technique for video inference named ShiftCaching, which can ensure temporal smoothness between video clips and reduce redundant computation.
  • We introduce and curate a new video virtual try-on dataset, TikTokDress, which encompasses a wide range of backgrounds and complex movements and features high-resolution videos, filling a gap that exists in previous video virtual try-on datasets.
Figure 2: Overview of Stage 2 of our SwiftTry framework. Stage 1 is similar except the input is single image frame and does not have temporal attention layers. Given an input video and a garment image, our method first extracts the masked video, corresponding masks, and pose sequence. The masked video is encoded into latent space by the VAE Encoder, which is then concatenated with noise, masks, and pose features before being processed by the Main U-Net. To inpaint the garment during the denoising process, we use a Garment U-Net and a CLIP encoder to extract both low- and high-level garment features. These features are integrated into the Main U-Net through spatial and cross-attention mechanisms.
Related Work
Image Virtual Try-On. Traditional image virtual try-on methods (Han et al. 2018; Wang et al. 2018a; Dong et al. 2019a; Yang et al. 2020; Ge et al. 2021; He, Song, and Xiang 2022; Choi et al. 2021; Lee et al. 2022; Xie et al. 2023) commonly employ a two-stage pipeline based on GANs (Goodfellow et al. 2014). In this approach, the target clothing is first warped and then fused with the person image to create the try-on effect. Various techniques have been utilized for clothing warping, including thin-plate spline (TPS) warping (Han et al. 2018), spatial transformer networks (STN) (Li et al. 2021), and flow estimation (Xie et al. 2023). Despite these advances, such methods often face limitations in generalization, resulting in significant performance degradation when applied to person images with complex backgrounds.
Recently, diffusion models have markedly enhanced the realism of images in generative tasks, leading to their increasing adoption in virtual try-on research. For instance, TryOnDiffusion (Zhu et al. 2023) presents a virtual try-on method utilizing two U-Nets, but it requires a large dataset of image pairs depicting the same person in various poses, which can be difficult to acquire. StableVITON (Kim et al. 2023) conditions the garment in a ControlNet (Zhang, Rao, and Agrawala 2023)-style using a zero cross-attention block, while IDM-VTON (Choi et al. 2024) proposes GarmentNet to encode low-level features combined with high-level semantic features extracted via IP-Adapter (Ye et al. 2023). Despite these advancements, extending these existing image virtual try-on methods for video often results in significant inter-frame inconsistency and flickering, which adversely affects the overall quality of the generated results.
Video Virtual Try-On. Several efforts have been made to develop virtual try-on for videos. FW-GAN (Dong et al. 2019b) integrates an optical flow prediction module from Video2Video (Wang et al. 2018b) to warp preceding frames to the current frame, enabling the synthesis of temporally coherent subsequent frames. MV-TON (Zhong et al. 2021) introduces a memory refinement module that retains and refines features from previous frames. ClothFormer (Jiang et al. 2022) utilizes a vision transformer in its try-on generator to minimize blurriness and temporal artifacts, and it features an innovative warping module that combines TPS-based and appearance-based methods to address issues like incorrect warping due to occlusions. Among diffusion-based methods, Tunnel Try-On (Xu et al. 2024b) is the first to apply diffusion models for video virtual try-on, effectively handling camera movement and maintaining consistency, although its demo videos are limited to only a few seconds long. ViViD (Fang et al. 2024) introduced a large-scale video try-on dataset with multiple categories, but it remains limited by simple backgrounds and movements, which constrains its ability to maintain long-term consistency and coherence. In this paper, we propose a novel technique that ensures temporal smoothness and coherence across video clips, complemented by a caching technique (Ma, Fang, and Wang 2024) that reduces redundant computations during long video inference.
Methods
Problem Statement: Given a source video $V=\{I_{1},I_{2},\dots,I_{N}\}\in\mathbb{R}^{N\times 3\times H\times W}$ of a person and a garment image $g\in\mathbb{R}^{3\times H\times W}$, where $N$, $H$, and $W$ represent the video length, frame height, and frame width, respectively, our goal is to synthesize an target video $\hat{V}=\{\hat{I}_{1},\hat{I}_{2},\dots,\hat{I}_{N}\}\in\mathbb{R}^{N\times 3\times H\times W}$ of the person wearing the garment, while preserving the motion of the person, the background in $V$, and the color and texture of $g$.
It is important to note that collecting both source and target videos of the same person with identical motion and gestures, differing only in the garment, is extremely challenging. As a result, most video try-on approaches adopt a self-supervised training method, where only a single video is used, and the garment regions are masked. The model is then trained to inpaint the masked regions using guidance from the garment image.
In the next section, we first describe our overall video try-on architecture and then discuss in detail the ShiftCaching technique – one of our main contributions.
Figure 3: Illustration of fully and partially computed frames with a chunk size $N=8$ and shift $\Delta=4$. In the partially computed chunk, one half uses cached features from timestep $t+2$, and the other half uses features from $t+1$.
Overall Architecture
Our approach consists of two stages: first, training a diffusion-based image try-on model, and then extending it to work with video data by incorporating temporal attention into every block of the Main UNet.
In the first stage, inspired by StableVITON (Kim et al. 2023), we design diffusion-based image try-on model with two submodules: the Garment UNet and the Main UNet, as illustrated in Fig. 2. The Main UNet is a modified inpainting model, initialized with the pretrained weights from Stable Diffusion (Rombach et al. 2022). It takes as input four channels of latent noise, four channels of latent representations of the masked image (i.e., the person image with the clothes region masked), and one channel for the binary mask representing the inpainting region. To further enhance the generation, we add the pose skeleton as an additional control, represented by a pose map rendered from DW-Pose (Yang et al. 2023). This results in a 13-channel input fed into the Main UNet, which predicts the cleaned latent over $T$ timesteps. Finally, the cleaned latent is passed through a decoder to produce the output image. The Garment UNet has a similar architecture to the Main UNet but only takes the garment image as its input, rather than the multiple channels used in the Main UNet. This module is designed to extract both detailed and high-level features from the garment, guiding the Main UNet to accurately replicate the garment’s appearance through Reference Attention. Specifically, we follow the Reference Attention from AnimateAnyone (Hu et al. 2023), replicating the Garment Unet’s feature maps along the temporal dimension and concatenating them with the Main Unet’s feature maps along the spatial dimension before undergoing Unet’s self-attention.
We use the VITON-HD dataset (Choi et al. 2021) to train our network during this stage.
[…]
$S=0$ 9.040 1.544
$S=4$ 8.822 1.176
$S=8$ 8.947 0.801
$S=15$ 8.675 0.104
[…]
Figure 4: Comparison of a fully computed frame and a partially computed frame. The partially computed frame uses Masked Temporal Attention instead of standard Temporal Attention to address mismatches in cached features.
In the second stage of designing our diffusion-based video try-on model, we ensure temporal consistency across video frames by adapting the Main UNet from the image try-on model. This adaptation involves converting its 2D layers into pseudo-3D layers (Guo et al. 2023; Wu et al. 2023; Zhou et al. 2022) and adding a temporal attention layer after the Spatial and Cross-Attention layers to capture temporal correlations between frames. The architecture of the modified UNet blocks is illustrated in Fig. 2. In the temporal attention layer, the features are reshaped into the shape of $(H\times W)\times N\times C$, where $C$ is the number of feature channels and $H\times W$ is the batch-size dimension, to compute self-attention along the temporal dimension. This allows a location in the latent space of frame $t$ to interact with the same location in other frames within a chunk of $N$ frames. This design is highly efficient as it avoids the expense of full 3D attention by factorizing it into two consecutive steps: spatial attention (2D) and temporal attention (1D). This approach allows a location in one frame to exchange information with every location in all other frames. Additionally, we incorporate sinusoidal positional encoding to help the model recognize the position of each frame in the video, following (Guo et al. 2023). In this stage, we train only the temporal attention layer while keeping the other layers unchanged, using a video dataset.
ShiftCaching Technique
Due to memory constraints, current video diffusion-based virtual try-on methods can only generate video chunks of 16 frames at a time. Previous approaches (Fang et al. 2024; He et al. 2024; Xu et al. 2024b) use a temporal aggregation technique (Tseng, Castellon, and Liu 2023; Xu et al. 2023) to stitch overlapping video clips into longer sequences. In this process, the long video is divided into overlapping clips with an overlap size $S$, typically set to $N/2$ or $N/4$. At each denoising timestep $t$, the overlapping noise predictions are merged using a simple averaging technique. However, this method involves a trade-off: a smaller overlap size, such as $S=4$, can cause temporal flickering and texture artifacts, while a larger overlap size, such as $S=15$, improves consistency but greatly slows down the process as shown in Tab. 1.
To achieve good temporal coherence and smoothness without recomputing the overlapped regions, we propose a shifting mechanism during inference. Specifically, we divide the long video into non-overlapping chunks ($S=0$). At each DDIM sampling timestep $t$, we shift these chunks by a predefined value $\Delta$ between two consecutive frames, allowing the model to process different compositions of noisy chunks at each step. An example of a fixed $\Delta=4$ applied to a chunk with length $N=8$ is illustrated in Fig. 3.
To further accelerate the inference process, we can skip a random chunk to reduce redundant computation during denoising. However, dropping chunks without adjustment can lead to abrupt changes in noise levels in the final results. Following (Ma, Fang, and Wang 2024), which notes that adjacent denoising steps share significant similarities in high-level features, we instead perform partial computations on the Main U-Net. Specifically, we use a cache to copy the latest features from the fully computed timestep $z_{t+1}$ (Red frame) and use these features to partially compute the current latent $z_{t}$ (White frame), bypassing the deeper blocks of the UNet, as illustrated in Fig. 4.
When performing partial computations on a chunk, the cached features typically include a first half from timestep $t+2$ and a second half from timestep $t+1$, which can cause mismatches between the two halves. To address this, we introduce a Masked Temporal Attention mechanism. This mechanism incorporates a special mask with size $N\times N$ during the softmax attention calculation to set the attention matrix values to 0, preventing information transfer from less accurate features (timestep $t+2$) to more accurate features (timestep $t+1$), while allowing transfer from good features to bad features. This mechanism ensures both smoothness and high quality in the partially computed cells.
TikTokDress Dataset
Label Gender Skin tone Camera position Distance Action Background
Male Female White Yellow Black Bottom Top Center Near Far Move Stay Dynamic Static
Counting 267 550 541 124 152 576 7 231 570 247 275 542 94 724
Table 2: Data statistics illustrating our dataset’s diversity and complexity. The table breaks down attributes such as gender, skin tone, camera positions, distances (Near/Far), and actions (Move/Stay), which indicate whether the actor is moving or stationary.
Figure 5: Example videos from the TikTokDress dataset highlighting diversity in skin tones, genders, camera angles, and clothing types.
Figure 6: SAM 2 (Ravi et al. 2024) failures due to sensitivity to prompts (green for positive, red for negative), with manual corrections needed for challenging areas (green rectangles).
Public datasets for single-image virtual try-ons, such as VITON-HD and DressCode, often suffer from simple backgrounds and limited human poses, despite offering high-quality images. These datasets are also restricted to single-image scenarios. Similarly, the VVT dataset, a standard for video virtual try-on, has notable drawbacks, including uniform movements, white backgrounds, and low resolution ($256\times 192$), making it unsuitable for real-world applications, particularly in the short-video industry where higher resolution is crucial. Given that real-world videos are typically recorded on mobile phones, which introduces variations in background, camera position, and lighting, there’s a need for a more robust dataset. To address these gaps, we introduce TikTokDress, a high-resolution video virtual try-on dataset that includes complex backgrounds, diverse movements, and gender representation. Each video is paired with the corresponding garment and is annotated with detailed human poses and precise binary cloth masks, enhancing its value for real-world applications.
First, the quality of garment masks in our dataset is crucial for enhancing try-on results, as shown in the Supplementary Material. While existing datasets like VITON-HD (Choi et al. 2021), DressCode (Morelli et al. 2022), and VVT (Dong et al. 2019b) use a standard segmenter (Li et al. 2020), it struggles with complex videos, resulting in subpar performance. In contrast, TikTokDress offers manually corrected, highly accurate garment masks, leading to significantly improved video try-on quality.
Second, our TikTokDress dataset captures a broad range of human poses and dynamic movements, such as dancing, common in short-form videos. As shown in Fig. 5, it includes variations in camera distance and diverse backgrounds, from indoor to outdoor settings with complex lighting. Additionally, it features a variety of clothing types, from casual T-shirts to structured garments like sweaters and challenging attire such as chainmail tops, addressing real-world challenges in video virtual try-on.
Video collection and annotation.
Our dataset consists of short TikTok clips (10-30 seconds) featuring various dance routines, as shown in Fig. 5. We expanded the TikTok Dataset (Jafarian and Park 2021) by adding videos to increase diversity in backgrounds, skin tones, and clothing styles, resulting in 817 video-garment pairs. Videos with excessive motion blur or low-quality garments were excluded. To ensure accurate garment matching, we manually selected high-quality matches from fashion retail websites. The dataset includes over 270,000 RGB frames extracted at 30 frames per second. We also calculated 2D keypoints and dense pose information using DWPose (Yang et al. 2023) and DensePose (Güler, Neverova, and Kokkinos 2018). Dataset statistics are summarized in Tab. 2, highlighting diversity in gender, skin tone, and camera positions.
Creating high-quality garment masks for each video was challenging due to the need for precise segmentation in every frame. We used SAM 2 (Ravi et al. 2024) to extract masks for both clothing and arms, but its sensitivity to prompt points and specific frames (see Fig. 6-(a)) required an additional solution. We developed an algorithm (described in the Supplemental Material) to automate the selection of optimal frames and prompts, improving efficiency across various scenarios. Despite these advancements, complex or unusual garments still needed manual intervention, which we detail with examples in the Supplementary Material. This meticulous process was essential for ensuring the dataset’s high quality and reliability.
Experiments
Datasets: We evaluate our approach on the VVT dataset (Dong et al. 2019b) and our new TikTokDress dataset. The VVT dataset, a standard benchmark for video virtual try-on, includes 791 paired videos of individuals and clothing images, with 661 for training and 130 for testing, all at $256\times 192$ resolution. The videos feature simple movements against plain backgrounds. In contrast, the TikTokDress dataset offers a more complex challenge, with varied backgrounds, dynamic movements, and diverse body poses. It comprises 693 training videos and 124 testing videos at $540\times 720$ resolution, totaling 232,843 frames for training and 39,705 frames for testing.
Metrics: We evaluate our approach using image-based and video-based metrics in both paired and unpaired settings, as outlined in (Jiang et al. 2022). In paired settings, we use SSIM (Wang et al. 2004) and LPIPS (Zhang et al. 2018) to assess reconstruction quality. In unpaired settings, we measure visual quality and temporal consistency with Video Fréchet Inception Distance (VFID) (Dong et al. 2019b). Additionally, we measure inference speed in frames per second (FPS) to demonstrate speed improvements.
Implementation details: The training process is divided into two stages. In the first stage, we focus on inpainting and preserving detailed garment textures using the VITON-HD dataset (Choi et al. 2021). We fine-tune the Garment UNet, Pose Encoder, and Main UNet decoder, initializing the Main UNet and Garment UNet with pretrained weights from SD 1.5, while keeping the VAE Encoder, Decoder, and CLIP image encoder weights unchanged. In the second stage, we incorporate temporal attention layers into the previously trained model, initializing these new modules with pretrained weights from AnimateDiff (Guo et al. 2023).
Method VVT
[…]
CP-VTON 0.535 0.459 6.361 12.100 N/A
FBAFN 0.157 0.870 4.516 8.690 N/A
StableVITON 0.184 0.760 17.068 11.254 0.241
StableVITON+AA 0.270 0.683 12.597 3.336 1.165
FWGAN 0.283 0.675 8.019 12.150 N/A
MVTON 0.068 0.853 8.367 9.702 N/A
ClothFormer 0.081 0.921 3.967 5.048 N/A
Tunnel Try-On 0.054 0.913 3.345 4.614 N/A
ViViD† 0.119 0.829 6.788 0.853 1.409
WildVidFit N/A N/A 4.202 N/A N/A
SwiftTry (ours) 0.066 0.887 3.289 0.534 2.270
Table 3: Comparisons on the VVT dataset (Dong et al. 2019b). † means our re-evaluation from the provided code.
Method TikTokDress
[…]
ViViD† 0.129 0.824 5.638 1.409
SwiftTry w/o ShiftCaching 0.075 0.891 3.865 1.177
SwiftTry (ours) 0.074 0.888 4.231 2.270
Table 4: Comparisons on the TikTokDress dataset.
[…]
FS 0.061 0.882 8.971 0.864 1.544
RS 0.060 0.883 8.878 0.853 1.544
FS, P 50% 0.060 0.883 8.932 0.887 2.270
RS, P 50% 0.060 0.883 8.938 0.888 2.270
Table 5: Study of our Temporal layers with ShiftCaching. FS: Fixed Shift, RS: Random Shift, P: Partially Computed.
[…]
FA 0.060 0.883 8.932 0.887
HA 0.059 0.886 8.679 0.796
QA 0.061 0.882 8.990 0.909
CA 0.086 0.854 13.520 5.501
Table 6: Ablation study of our Temporal layers with ShiftCaching. FA: Full Attention, HA: Half Attention, QA: Quarter Attention, CA: Causal Attention.
[…]
8 0.062 0.880 9.312 1.027 0.914
16 0.061 0.881 8.822 0.851 1.176
24 0.163 0.821 8.724 0.800 1.723
Table 7: Ablation study of testing with 8, 16, and 24 frames, with the default training set to 16 frames.
Comparisons with Prior Approaches
We compare our approach with other video virtual try-on methods using the VVT and TikTokDress datasets. As most methods are closed-source, we rely on reported results and available generated videos for comparison. For GAN-based methods, we evaluate against FW-GAN (Dong et al. 2019b), MV-TON (Zhong et al. 2021), and ClothFormer (Jiang et al. 2022). For diffusion-based methods, we compare with Tunnel Try-On (Xu et al. 2024b), ViViD (Fang et al. 2024), and WildVidFit (He et al. 2024). We re-evaluate ViViD (Fang et al. 2024) on the VVT dataset, as it is the only method with available inference code and pre-trained weights. Additionally, we compare our model with the image-based virtual try-on method StableVITON (Kim et al. 2023), finetuned on the VVT dataset, in a frame-by-frame manner. We also evaluate a baseline combining StableVITON and AnimateAnyone (Hu et al. 2023), where StableVITON performs the try-on for individual frames, and AnimateAnyone generates a video based on the source motion.
Figure 7: Qualitative results of our method on the TikTokDress dataset.
Figure 8: Qualitative comparison with prior method on the VVT dataset.
Quantitative results: Tab. 3 presents the comparison on the VVT dataset. Our approach excels in the VFID metric, indicating superior visual quality and consistency, and also performs competitively in SSIM and LPIPS scores. While ClothFormer achieves a high SSIM score, its VFID is lower due to the limitations of its GAN-based method. Our ShiftCaching technique enhances performance, increasing the frame rate to 2.27 FPS – over 1.5 times faster. We also evaluated our method on the TikTokDress dataset, as detailed in Tab. 4. Our analysis shows that while these methods produce accurate individual frames, they often struggle with flickering and inconsistencies due to poor temporal coherence and motion handling across frames.
Qualitative results: As shown in Fig. 8 and Fig. 7, the textures on the garment vary between frames. Additionally, there are significant jitters between adjacent frames with these methods, which can be observed more intuitively in videos provided in our Supplementary Material.
Ablation Study
We conducted ablation studies on the VVT dataset to investigate various factors affecting the performance of SwiftTry.
Study on the ShiftCaching Technique is shown in Tab. 5. The results indicate that using random shifts provides the best consistency. When combined with partial computation of 50% of the frames, this approach accelerates inference by 1.5 times while maintaining comparable quantitative metrics to other methods.
Study on Different Types of Masks in Masked Temporal Attention is shown in Tab. 6. The results reveal that Half Attention yields the best performance. This suggests that allowing only the bad features (from timestep $t+2$) to access the good features (from timestep $t+1$) and allowing only the good features to interact with each other, produces the optimal results. Detailed explanations of different masking attention are described in our Supplementary Material.
Impact of Inference Video Chunk Length is examined in Tab. 7. The study reveals that matching the training and inference video chunk lengths – both set to $N=16$ – yields the best results.
Conclusion
In conclusion, we have proposed a novel technique, ShiftCaching, which ensures temporal smoothness across video clips while effectively reducing redundant computations during video inference. This advancement enhances the efficiency and quality of video virtual try-on, making it more practical for real-world applications. Additionally, we have introduced a new dataset, TikTokDress, designed specifically for video virtual try-on. This dataset stands out for its diverse range of backgrounds, complex movements, and high-resolution videos, addressing the limitations of existing datasets and providing a valuable resource for future research in this area.
Supplementary Material
In this supplementary material, we present additional experimental results and details that could not be included in the main paper due to space constraints. First, we offer a comprehensive overview of the training process for our proposed method. Next, we delve into the specifics of various masking strategies employed in the attention module. We also provide detailed information of creating our TikTokDress dataset. Finally, we showcase additional qualitative results of our method on the VVT and TikTokDress datasets.
Appendix A Training Detail
Optimization. We use Adam (Kingma and Ba 2014) optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, a fixed learning rate of $51e-$ 5 for both image-pretraining stage and stage 2. We also utilize Min-SNR Weighting Strategy (Hang et al. 2023) with $\gamma=5.0$. We conduct our training in mixed-precision fp16.
Hardware. All the codes are implemented in Pytorch utilizing Diffusers (von Platen et al. 2022) framework and we only conduct training on a single A100 40GB in both two stages. The training takes around 3 days for stage 1 on VITON-HD (Choi et al. 2021) and around 30 hours for stage 2 on both VVT and TikTokDress. Additionally, the FPS benchmarks reported in the paper are also obtained using the A100 GPU.
Training. In the first stage, we optimize the Main Unet decoder, Garment Unet, and Pose Encoder on VITON-HD (Choi et al. 2021) while other modules are kept fixed. The training resolution at this stage remains at $1024\times 768$, consistent with the original resolution. Then, we sample a clip of 16 frames with resolution $512\times 384$ for each video as input for stage 2. We optimize motion modules only in this stage.
Appendix B Different Masking Attention Mechanisms
Fig. 9 illustrates the different attention masking strategies in our ShiftCaching method. In the Full Attention configuration ($a$), the attention mask is filled with $0.0$, allowing each frame to attend to all others. However, this can lead to less accurate features (i.e., $t+2$, shown in white) influencing the more accurate ones (i.e., $t+1$, shown in red). In the Half Attention ($b$) setup, the query frame always attends to the accurate features ($t+1$), enabling the transfer of quality features to less accurate ones. The Quarter Attention($c$) strategy allows frames with less accurate features to attend to all other frames while ensuring that the accurate features do not attend to less accurate ones, thereby preserving their quality. Finally, Causal Attention ($d$) ensures that the computation for each frame should only depend on the succeeding frames. We perform experiments and find out that Half Attention($b$) yields the best performance.
Figure 9: The attention mask used in different masked temporal attention mechanisms in ShiftCaching. White cells in the matrix represent a value of 0.0, while gray cells indicate a value of $\infty$- ∞.
Appendix C TikTokDress Dataset
Fig. 12 showcases examples from our dataset, which is diverse in background (ranging from indoor to outdoor settings), camera positions (from top to bottom angles), person distances (from far to close to the camera), skin tones (including black, yellow, and white), and garment types (such as t-shirts, sweaters, crop tops, long-sleeve t-shirts, and dresses). Based on this diversity, we believe that our dataset better reflects real-life scenarios for virtual try-ons.
Figure 10: Impact of mask quality on try-on results. The first row shows target clothing images. The second row demonstrates the effects of an inaccurate mask, leading to flawed try-on results in the third row. The fourth row highlights how a precise mask significantly improves quality and accuracy. Accurate masks are crucial for high-quality try-ons.
Figure 11: Examples of T-pose
Agnostic masks are crucial for achieving accurate try-on results. For example, in Fig. 10, an incorrect agnostic mask (highlighted by the red circle) causes the model to hallucinate a garment to fit the mask, resulting in a try-on outcome where the garment appears as a tank top (highlighted by the blue rectangle in the first row). With a correct garment mask, a good try-on result is achieved (highlighted by the green rectangle in the first row). In the second row, a garment that is not fully masked leads to the model hallucinating additional red stripes in the middle of the red garment (highlighted by the blue rectangle in the second row). Again, a correct garment mask produces a good try-on result (highlighted by the green rectangle in the second row).
Figure 12: Examples from our dataset, which feature diverse poses and high-quality garments.
To obtain accurate agnostic masks, we use SAM2 (Ravi et al. 2024) for both garment and arm segmentation. Given a video with multiple frames showing diverse poses, we select the frame that displays as many visible body parts as possible and is clear, similar to Fig. 11. We apply the algorithm described in Algorithm 1 for this frame selection.
The agnostic masks are a combination of the garment mask and the arm mask. For garment masks, we choose positive points in the prompting process, specifically the body center, right shoulder, and left shoulder. The body center is defined as the midpoint between the neck, right hip, and left hip (see points 1, 8, and 11 in Fig. 11). For the negative points, we select noise, the left eye, and the right eye. For arm masks, the positive points we choose are the left elbow, right elbow, left wrist, and right wrist. For the negative points, we choose the right knee, nose, and left knee. After obtaining the garment and arm masks, we combine them and dilate the result three to seven times to ensure it fully conceals the original garment’s shape.
Garment Image. To extract garment images from videos, we use the Google Lens engine to find the best match at the highest possible resolution. In cases where Google Lens does not return a satisfactory result, we manually search on Google using keywords related to the type of clothing, color, and brand, selecting the best match. If we are unable to find a suitable garment image, we exclude that video from our dataset. After obtaining the garment image, in some cases, we horizontally flip the image to align it with the video.
Algorithm 1 Selecting the Optimal Frame for SAM2 Prompting
1:  Define:
2:      kps $\leftarrow$ keypoints
3:      p_joints $\leftarrow$ pre_define_joints
4:      v_p $\leftarrow$ visible_part
5:  Initialize:
[…]
8:      frame_kps_scores $\leftarrow$ []
9:  Function perfect_pose_score(keypoints):
10:      score $\leftarrow$ 0
11:      visible_part $\leftarrow$ 0
12:      for each key in p_joints do:
13:         j_A, j_B, j_C $\leftarrow$ p_joints[key]
14:         angle $\leftarrow$ calculate_angle(kps[j_A], kps[j_B], kps[j_C])
[…]
16:            v_p $\leftarrow$ v_p + 1
17:      return score, v_p
18:  Iterate through frames:
19:      for frame_idx, kps in frames do:
20:         score, v_p $\leftarrow$ perfect_pose_score(kps)
21:         frame_kps_scores.append((score, frame_idx, v_p))
22:  Sort frames by visible parts and score:
23:      frame_kps_scores.sort(key=lambda x: (-x[2], x[0]))
24:  Select the frame with the highest visible parts and perfect body pose score.
Appendix D More Visualization Results
Qualitative results on the VVT dataset
In the video try-on task, we compare our method on the VVT dataset (Dong et al. 2019b) against ViViD (Fang et al. 2024) in Fig. 13. As demonstrated in video $b.mp4$(see in Supplemental Videos), our method consistently preserves the garment texture throughout the entire video better than ViViD in maintaining texture stability over time. In video $d.mp4$ (see in Supplemental Videos), the “GAP” letters on the garment in ViViD exhibit noticeable flickering, a result of its temporal averaging technique. In contrast, our method, which employs the ShiftCaching technique, significantly reduces this flickering, ensuring smoother and more stable visual results (Notes: we recommend to see in the supplementary videos).
Figure 13: Qualitative comparison between our method and ViViD. Our model achieved robust results through long video, thereby forming a coherent video sequence.
Figure 14: More examples of our video try-on results on our TikTokDress dataset. The first row illustrates how well our method’s performs in various backgrounds. The second row highlights our method’s ability to preserve texture consistency over time.
Qualitative results on the TikTokDress dataset.
Fig. 14 showcases additional results of our video virtual try-on method on the TikTokDress dataset. By utilizing SAM 2 (Ravi et al. 2024) and manually refining masks on our dataset, combining with the SAM2 frame selection algorithm Algorithm 1, our try-on results demonstrate superior performance compared to those trained on the VVT dataset. Our method ensures consistent texture quality for garments and, when combined with ShiftCaching, effectively produces continuous and smooth long videos (Notes: we recommend to see more detail in the supplementary videos).

### v2

@@
 Given an input video of a person and a new garment, the objective of this paper is to synthesize a new video where the person is wearing the specified garment while maintaining spatiotemporal consistency.
-While significant advances have been made in image-based virtual try-ons, extending these successes to video often results in frame-to-frame inconsistencies.
-Some approaches have attempted to address this by increasing the overlap of frames across multiple video chunks, but this comes at a steep computational cost due to the repeated processing of the same frames, especially for long video sequence.
-To address these challenges, we reconceptualize video virtual try-on as a conditional video inpainting task, with garments serving as input conditions.
+Although significant advances have been made in image-based virtual try-on, extending these successes to video often leads to frame-to-frame inconsistencies.
+Some approaches have attempted to address this by increasing the overlap of frames across multiple video chunks, but this comes at a steep computational cost due to the repeated processing of the same frames, especially for long video sequences.
+To tackle these challenges, we reconceptualize video virtual try-on as a conditional video inpainting task, with garments serving as input conditions.
 Specifically, our approach enhances image diffusion models by incorporating temporal attention layers to improve temporal coherence.
-To reduce computational overhead, we introduce ShiftCaching, a novel technique that maintains temporal consistency while minimizing redundant computations.
+To reduce computational overhead, we propose ShiftCaching, a novel technique that maintains temporal consistency while minimizing redundant computations.
 Furthermore, we introduce the TikTokDress dataset, a new video try-on dataset featuring more complex backgrounds, challenging movements, and higher resolution compared to existing public datasets.
-Extensive experiments show that our approach outperforms current baselines, particularly in terms of video consistency and inference speed.
-Data and code are available at https://github.com/VinAIResearch/swift-try.
+Extensive experiments demonstrate that our approach outperforms current baselines, particularly in terms of video consistency and inference speed.
+The project page is available at https://swift-try.github.io/.
 Introduction
-Figure 1: Results of our SwiftTry compared with those of ViViD (Fang et al.
-2024), the previous method for video try-on.
-Our method maintains detail and consistency in the texture of the garments and runs over 60% faster.
+Figure 1: Results of our SwiftTry compared to those of ViViD (Fang et al.
+2024), a previous method for video try-on.
+Our method preserves garment texture detail and consistency while achieving over 60% faster runtime.
 Video virtual try-on is an emerging research area (Chen et al.
@@
 2024) with significant potential in fashion and e-commerce.
-The ability to realistically visualize how a garment looks on a person in a video could transform online shopping.
-However, despite recent progress in image-based virtual try-on (He, Song, and Xiang 2022; Choi et al.
+The ability to realistically visualize how a garment appears on a person in a video could transform online shopping.
+However, despite recent advances in image-based virtual try-on (He, Song, and Xiang 2022; Choi et al.
 2021; Lee et al.
@@
 2024; Fang et al.
-2024) often struggle with temporal inconsistencies, leading to visual artifacts and flickering between frames, which undermines the realism of the virtual try-on experience.
-Additionally, the high computational cost of rendering high-quality results over extended sequences hampers the practicality of these approaches for real-world applications.
-Another challenge is lacking an evaluation dataset.
-The first public video try-on dataset VVT (Dong et al.
-2019b) only covers basic pattern garments, form-fitting T-shirts, uniform backgrounds, static camera angles, and repetitive human motions.
-Recently, ViViD (Fang et al.
-2024) released the first practical dataset for video virtual try-on; however, it struggles to handle in-the-wild scenarios, such as complex movements and diverse backgrounds, making it challenging to meet the demands of practical applications.
-Moreover, one reason for the poor quality video try-on result is primarily due to masks extracted using human parsing segmentation (Li et al.
-2020) applied on each frame of the video.
-In this paper, we tackle these challenges by introducing two key contributions.
-First, we present a new high-quality dataset named TikTokDress consisting of 817 videos which is specifically designed for training and evaluating video virtual try-on models.
+2024) often struggle with temporal inconsistencies, resulting in visual artifacts and flickering between frames, which undermines the realism of the virtual try-on experience.
+Additionally, the high computational cost of rendering high-quality results over extended sequences limits the practicality of these approaches in real-world applications.
+Another challenge is the lack of an adequate evaluation dataset.
+The first public video try-on dataset, VVT (Dong et al.
+2019b), only includes basic pattern garments, form-fitting T-shirts, uniform backgrounds, static camera angles, and repetitive human motions.
+More recently, ViViD (Fang et al.
+2024) introduced the first practical dataset for video virtual try-on.
+However, it struggles to handle in-the-wild scenarios, such as complex movements and diverse backgrounds, making it difficult to meet the demands of real-world applications.
+Moreover, the poor quality of video try-on results can often be attributed to the inaccurate masks extracted using human parsing segmentation (Li et al.
+2020), which are applied to each frame of the video.
+In this paper, we address these challenges with two key contributions.
+First, we introduce a new high-quality dataset, named TikTokDress, consisting of 817 videos specifically designed for training and evaluating video virtual try-on models.
 This dataset features realistic scenes, diverse garment types, and complex movements, providing a robust foundation for advancing research in this field.
-Second, we introduce a novel video virtual try-on framework named SwiftTry, as shown in Fig.
-1 which significantly reduces the computational cost of processing long video sequences while maintaining temporal consistency.
+Second, we propose a novel video virtual try-on framework named SwiftTry, as illustrated in Fig.
+1, which significantly reduces the computational cost of processing long video sequences while maintaining temporal consistency.
 Our framework is inspired by state-of-the-art diffusion-based image virtual try-on methods (Kim et al.
@@
 2024a; Choi et al.
-2024) and incorporates temporal attention in the UNet architecture to train on video try-on data.
-During inference, we introduce a new technique called ShiftCaching, which ensures temporal coherence and smooth transitions between video clips while also reducing redundant computation compared to previous methods.
-Extensive experimental results indicate that our proposed SwiftTry framework, incorporating the aforementioned techniques, substantially surpasses other video virtual try-on methods in performance.
+2024) and incorporates temporal attention within the UNet architecture to train on video try-on data.
+During inference, we introduce a new technique called ShiftCaching, which ensures temporal coherence and smooth transitions between video chunks while minimizing redundant computation compared to previous methods.
+Extensive experimental results demonstrate that our proposed SwiftTry framework, leveraging these techniques, significantly outperforms existing video virtual try-on methods in both accuracy and efficiency.
 In summary, the contributions of our work are as follows:
-  • We propose a new technique for video inference named ShiftCaching, which can ensure temporal smoothness between video clips and reduce redundant computation.
+  • We propose a new technique for video inference named ShiftCaching, which can ensure temporal smoothness between video chunks and reduce redundant computation.
   • We introduce and curate a new video virtual try-on dataset, TikTokDress, which encompasses a wide range of backgrounds and complex movements and features high-resolution videos, filling a gap that exists in previous video virtual try-on datasets.
-Figure 2: Overview of Stage 2 of our SwiftTry framework.
-Stage 1 is similar except the input is single image frame and does not have temporal attention layers.
+Figure 2: Overview of Stage 2 of our SwiftTry framework (Note that stage 1 is similar, except the input is a single image frame, and it does not include temporal attention layers).
 Given an input video and a garment image, our method first extracts the masked video, corresponding masks, and pose sequence.
-The masked video is encoded into latent space by the VAE Encoder, which is then concatenated with noise, masks, and pose features before being processed by the Main U-Net.
+The masked video is encoded into the latent space by the VAE Encoder, which is then concatenated with noise, masks, and pose features before being processed by the Main U-Net.
 To inpaint the garment during the denoising process, we use a Garment U-Net and a CLIP encoder to extract both low- and high-level garment features.
@@
 For instance, TryOnDiffusion (Zhu et al.
-2023) presents a virtual try-on method utilizing two U-Nets, but it requires a large dataset of image pairs depicting the same person in various poses, which can be difficult to acquire.
+2023) presents a virtual try-on method utilizing two U-Nets, but it requires a large dataset of image pairs of the same person in different poses, which can be difficult to acquire.
 StableVITON (Kim et al.
@@
 Video Virtual Try-On.
-Several efforts have been made to develop virtual try-on for videos.
+Several efforts have been made to develop virtual try-on systems for videos.
 FW-GAN (Dong et al.
-2019b) integrates an optical flow prediction module from Video2Video (Wang et al.
+2019b) incorporates an optical flow prediction module from Video2Video (Wang et al.
 2018b) to warp preceding frames to the current frame, enabling the synthesis of temporally coherent subsequent frames.
@@
 ClothFormer (Jiang et al.
-2022) utilizes a vision transformer in its try-on generator to minimize blurriness and temporal artifacts, and it features an innovative warping module that combines TPS-based and appearance-based methods to address issues like incorrect warping due to occlusions.
+2022) employs a vision transformer in its try-on generator to minimize blurriness and temporal artifacts.
+It also features an innovative warping module that combines TPS-based and appearance-based methods to address challenges such as incorrect warping caused by occlusions.
 Among diffusion-based methods, Tunnel Try-On (Xu et al.
-2024b) is the first to apply diffusion models for video virtual try-on, effectively handling camera movement and maintaining consistency, although its demo videos are limited to only a few seconds long.
+2024b) is the first to apply diffusion models for video virtual try-on, effectively handling camera movement and maintaining consistency.
+However, its demo videos are limited to only a few seconds in length.
 ViViD (Fang et al.
-2024) introduced a large-scale video try-on dataset with multiple categories, but it remains limited by simple backgrounds and movements, which constrains its ability to maintain long-term consistency and coherence.
-In this paper, we propose a novel technique that ensures temporal smoothness and coherence across video clips, complemented by a caching technique (Ma, Fang, and Wang 2024) that reduces redundant computations during long video inference.
+2024) introduced a large-scale video try-on dataset with multiple categories, but it remains limited by simple backgrounds and movements, which constrain its ability to ensure long-term consistency and coherence.
+In this paper, we propose a novel approach that establishes temporal smoothness and coherence across video chunks.
+Additionally, we integrate a caching technique (Ma, Fang, and Wang 2024) to reduce redundant computations during long video inference, significantly improving efficiency.
 Methods
-Problem Statement: Given a source video $V=\{I_{1},I_{2},\dots,I_{N}\}\in\mathbb{R}^{N\times 3\times H\times W}$ of a person and a garment image $g\in\mathbb{R}^{3\times H\times W}$, where $N$, $H$, and $W$ represent the video length, frame height, and frame width, respectively, our goal is to synthesize an target video $\hat{V}=\{\hat{I}_{1},\hat{I}_{2},\dots,\hat{I}_{N}\}\in\mathbb{R}^{N\times 3\times H\times W}$ of the person wearing the garment, while preserving the motion of the person, the background in $V$, and the color and texture of $g$.
-It is important to note that collecting both source and target videos of the same person with identical motion and gestures, differing only in the garment, is extremely challenging.
+Problem Statement: Given a source video $V=\{I_{1},I_{2},\dots,I_{N}\}\in\mathbb{R}^{N\times 3\times H\times W}$ of a person and a garment image $g\in\mathbb{R}^{3\times H\times W}$, where $N$, $H$, and $W$ represent the video length, frame height, and frame width, respectively, our goal is to synthesize a target video $\hat{V}=\{\hat{I}_{1},\hat{I}_{2},\dots,\hat{I}_{N}\}\in\mathbb{R}^{N\times 3\times H\times W}$ of the person wearing the garment, while preserving the motion of the person, the background in $V$, and the color and texture of $g$.
+It is important to note that collecting both source and target videos of the same person with identical motions and gestures, differing only in the garment, is extremely challenging.
 As a result, most video try-on approaches adopt a self-supervised training method, where only a single video is used, and the garment regions are masked.
@@
 In the next section, we first describe our overall video try-on architecture and then discuss in detail the ShiftCaching technique – one of our main contributions.
-Figure 3: Illustration of fully and partially computed frames with a chunk size $N=8$ and shift $\Delta=4$.
-In the partially computed chunk, one half uses cached features from timestep $t+2$, and the other half uses features from $t+1$.
+Figure 3: Illustration of fully and partially computed frames with a chunk size of $N=8$ and a shift of $\Delta=4$.
+In the partially computed chunk, one half uses cached features from $t+2$, while the other half uses features from $t+1$.
 Overall Architecture
-Our approach consists of two stages: first, training a diffusion-based image try-on model, and then extending it to work with video data by incorporating temporal attention into every block of the Main UNet.
+Our approach comprises two stages: first, training a diffusion-based image try-on model, and second, extending it to work with video data by incorporating temporal attention into every block of the Main UNet.
 In the first stage, inspired by StableVITON (Kim et al.
-2023), we design diffusion-based image try-on model with two submodules: the Garment UNet and the Main UNet, as illustrated in Fig.
+2023), we design a diffusion-based image try-on model with two submodules: the Garment UNet and the Main UNet, as illustrated in Fig.
 2.
-The Main UNet is a modified inpainting model, initialized with the pretrained weights from Stable Diffusion (Rombach et al.
+The Main UNet is a modified inpainting model initialized with pretrained weights from Stable Diffusion (Rombach et al.
 2022).
-It takes as input four channels of latent noise, four channels of latent representations of the masked image (i.e., the person image with the clothes region masked), and one channel for the binary mask representing the inpainting region.
-To further enhance the generation, we add the pose skeleton as an additional control, represented by a pose map rendered from DW-Pose (Yang et al.
+It takes as input four channels of latent noise, four channels of latent representations of the masked image (i.e., the person image with the clothing region masked), and one channel for the binary mask representing the inpainting region.
+To further enhance generation quality, we add the pose skeleton as an additional control, represented by a pose map rendered from DW-Pose (Yang et al.
 2023).
-This results in a 13-channel input fed into the Main UNet, which predicts the cleaned latent over $T$ timesteps.
+This results in a 13-channel input, which is fed into the Main UNet to predict the cleaned latent over $T$ timesteps.
 Finally, the cleaned latent is passed through a decoder to produce the output image.
@@
 This module is designed to extract both detailed and high-level features from the garment, guiding the Main UNet to accurately replicate the garment’s appearance through Reference Attention.
-Specifically, we follow the Reference Attention from AnimateAnyone (Hu et al.
-2023), replicating the Garment Unet’s feature maps along the temporal dimension and concatenating them with the Main Unet’s feature maps along the spatial dimension before undergoing Unet’s self-attention.
-We use the VITON-HD dataset (Choi et al. 2021) to train our network during this stage.
+Specifically, we follow the Reference Attention mechanism from AnimateAnyone (Hu et al.
+2023), replicating the Garment UNet’s feature maps along the temporal dimension and concatenating them with the Main UNet’s feature maps along the spatial dimension before applying the UNet’s self-attention.
+We use the VITON-HD dataset (Choi et al.
+2021) to train our network during this stage.
 […]
@@
 […]
-Figure 4: Comparison of a fully computed frame and a partially computed frame.
-The partially computed frame uses Masked Temporal Attention instead of standard Temporal Attention to address mismatches in cached features.
-In the second stage of designing our diffusion-based video try-on model, we ensure temporal consistency across video frames by adapting the Main UNet from the image try-on model.
+Figure 4: Comparison between a fully computed frame and a partially computed frame.
+The partially computed frame employs Masked Temporal Attention instead of standard Temporal Attention to resolve mismatches in cached features.
+In the second stage, we modify the Main UNet from the image try-on model to ensure temporal consistency across video frames.
 This adaptation involves converting its 2D layers into pseudo-3D layers (Guo et al.
@@
 2024b) use a temporal aggregation technique (Tseng, Castellon, and Liu 2023; Xu et al.
-2023) to stitch overlapping video clips into longer sequences.
-In this process, the long video is divided into overlapping clips with an overlap size $S$, typically set to $N/2$ or $N/4$.
+2023) to stitch overlapping video chunks into longer sequences.
+In this process, the long video is divided into overlapping chunks with an overlap size $S$, typically set to $N/2$ or $N/4$.
 At each denoising timestep $t$, the overlapping noise predictions are merged using a simple averaging technique.
@@
 To further accelerate the inference process, we can skip a random chunk to reduce redundant computation during denoising.
-However, dropping chunks without adjustment can lead to abrupt changes in noise levels in the final results.
+However, naively dropping chunks without adjustment can lead to abrupt changes in noise levels in the final results.
 Following (Ma, Fang, and Wang 2024), which notes that adjacent denoising steps share significant similarities in high-level features, we instead perform partial computations on the Main U-Net.
@@
 4.
-When performing partial computations on a chunk, the cached features typically include a first half from timestep $t+2$ and a second half from timestep $t+1$, which can cause mismatches between the two halves.
+When performing partial computations on a chunk, the cached features typically include the first half from timestep $t+2$ and the second half from timestep $t+1$, which can lead to mismatches between the two halves.
 To address this, we introduce a Masked Temporal Attention mechanism.
-This mechanism incorporates a special mask with size $N\times N$ during the softmax attention calculation to set the attention matrix values to 0, preventing information transfer from less accurate features (timestep $t+2$) to more accurate features (timestep $t+1$), while allowing transfer from good features to bad features.
-This mechanism ensures both smoothness and high quality in the partially computed cells.
+This mechanism applies a special mask of size $N\times N$ during the softmax attention calculation to set specific values in the attention matrix to 00.
+This prevents the transfer of information from less accurate features (timestep $t+2$) to more accurate features (timestep $t+1$), while allowing transfer from good features to bad features.
+This approach ensures both smoothness and high quality in the partially computed cells.
 TikTokDress Dataset
 Label Gender Skin tone Camera position Distance Action Background
-Male Female White Yellow Black Bottom Top Center Near Far Move Stay Dynamic Static
+Male Female White Asian Black Bottom Top Center Near Far Move Stay Dynamic Static
 Counting 267 550 541 124 152 576 7 231 570 247 275 542 94 724
-Table 2: Data statistics illustrating our dataset’s diversity and complexity.
-The table breaks down attributes such as gender, skin tone, camera positions, distances (Near/Far), and actions (Move/Stay), which indicate whether the actor is moving or stationary.
+Table 2: Data statistics highlighting the diversity and complexity of our dataset.
+The table provides a breakdown of attributes such as gender, skin tone, camera positions, distances (Near/Far), and actions (Move/Stay), indicating whether the actor is moving or stationary
 Figure 5: Example videos from the TikTokDress dataset highlighting diversity in skin tones, genders, camera angles, and clothing types.
-Figure 6: SAM 2 (Ravi et al. 2024) failures due to sensitivity to prompts (green for positive, red for negative), with manual corrections needed for challenging areas (green rectangles).
+Figure 6: SAM 2 failures due to its sensitivity to prompts (green for positive, red for negative), requiring manual corrections for challenging areas (green rectangles).
 Public datasets for single-image virtual try-ons, such as VITON-HD and DressCode, often suffer from simple backgrounds and limited human poses, despite offering high-quality images.
@@
 Similarly, the VVT dataset, a standard for video virtual try-on, has notable drawbacks, including uniform movements, white backgrounds, and low resolution ($256\times 192$), making it unsuitable for real-world applications, particularly in the short-video industry where higher resolution is crucial.
-Given that real-world videos are typically recorded on mobile phones, which introduces variations in background, camera position, and lighting, there’s a need for a more robust dataset.
-To address these gaps, we introduce TikTokDress, a high-resolution video virtual try-on dataset that includes complex backgrounds, diverse movements, and gender representation.
-Each video is paired with the corresponding garment and is annotated with detailed human poses and precise binary cloth masks, enhancing its value for real-world applications.
+Since real-world videos are typically recorded on mobile phones, which introduce variations in background, camera position, and lighting, there is a pressing need for a more robust dataset.
+To address these shortcomings, we introduce TikTokDress, a high-resolution video virtual try-on dataset that includes complex backgrounds, diverse movements, and balanced gender representation.
+Each video is paired with its corresponding garment and annotated with detailed human poses and precise binary cloth masks, enhancing its utility for real-world applications.
 First, the quality of garment masks in our dataset is crucial for enhancing try-on results, as shown in the Supplementary Material.
@@
 Video collection and annotation.
-Our dataset consists of short TikTok clips (10-30 seconds) featuring various dance routines, as shown in Fig.
+Our dataset consists of short TikTok clips (10–30 seconds) showcasing various dance routines, as illustrated in Fig.
 5.
-We expanded the TikTok Dataset (Jafarian and Park 2021) by adding videos to increase diversity in backgrounds, skin tones, and clothing styles, resulting in 817 video-garment pairs.
+We expanded the TikTok Dataset (Jafarian and Park 2021) by adding videos to enhance diversity in backgrounds, skin tones, and clothing styles, resulting in 817 video-garment pairs.
 Videos with excessive motion blur or low-quality garments were excluded.
-To ensure accurate garment matching, we manually selected high-quality matches from fashion retail websites.
+To ensure accurate garment matching, we manually curated high-quality matches from fashion retail websites.
 The dataset includes over 270,000 RGB frames extracted at 30 frames per second.
-We also calculated 2D keypoints and dense pose information using DWPose (Yang et al.
+Additionally, we computed 2D keypoints and dense pose information using DWPose (Yang et al.
 2023) and DensePose (Güler, Neverova, and Kokkinos 2018).
-Dataset statistics are summarized in Tab.
-2, highlighting diversity in gender, skin tone, and camera positions.
-Creating high-quality garment masks for each video was challenging due to the need for precise segmentation in every frame.
+Dataset statistics, summarized in Tab.
+2, highlight its diversity in gender, skin tone, and camera positions.
+Creating high-quality garment masks for each video posed a significant challenge due to the need for precise segmentation in every frame.
 We used SAM 2 (Ravi et al.
-2024) to extract masks for both clothing and arms, but its sensitivity to prompt points and specific frames (see Fig.
-6-(a)) required an additional solution.
-We developed an algorithm (described in the Supplemental Material) to automate the selection of optimal frames and prompts, improving efficiency across various scenarios.
-Despite these advancements, complex or unusual garments still needed manual intervention, which we detail with examples in the Supplementary Material.
+2024) to extract masks for both clothing and arms.
+However, its sensitivity to prompt points and specific frames (see Fig.
+6-(a)) necessitated an additional solution.
+To improve efficiency, we developed an algorithm (detailed in the Supplemental Material) for optimal frame and prompt selection.
+Complex garments still required manual refinement, as shown in the Supplementary Material.
 This meticulous process was essential for ensuring the dataset’s high quality and reliability.
@@
 WildVidFit N/A N/A 4.202 N/A N/A
-SwiftTry (ours) 0.066 0.887 3.289 0.534 2.270
+SwiftTry (ours) 0.066 0.887 3.589 0.534 2.270
 Table 3: Comparisons on the VVT dataset (Dong et al. 2019b). † means our re-evaluation from the provided code.
@@
 Conclusion
-In conclusion, we have proposed a novel technique, ShiftCaching, which ensures temporal smoothness across video clips while effectively reducing redundant computations during video inference.
+In conclusion, we have proposed a novel technique, ShiftCaching, which ensures temporal smoothness across video chunks while effectively reducing redundant computations during video inference.
 This advancement enhances the efficiency and quality of video virtual try-on, making it more practical for real-world applications.
@@
 In this supplementary material, we present additional experimental results and details that could not be included in the main paper due to space constraints.
-First, we offer a comprehensive overview of the training process for our proposed method.
-Next, we delve into the specifics of various masking strategies employed in the attention module.
-We also provide detailed information of creating our TikTokDress dataset.
-Finally, we showcase additional qualitative results of our method on the VVT and TikTokDress datasets.
+First, we provide a comprehensive overview of the training process for our proposed method.
+Next, we elaborate on the various masking strategies employed in the attention module.
+We also include detailed information about the creation of our TikTokDress dataset.
+Finally, we showcase additional qualitative results of our method on both the VVT and TikTokDress datasets.
 Appendix A Training Detail
 Optimization.
-We use Adam (Kingma and Ba 2014) optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, a fixed learning rate of $51e-$ 5 for both image-pretraining stage and stage 2.
-We also utilize Min-SNR Weighting Strategy (Hang et al.
+We use the Adam (Kingma and Ba 2014) optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, and a fixed learning rate of $51e-$ 5 for both the image pretraining stage and stage 2.
+We also adopt the Min-SNR Weighting Strategy (Hang et al.
 2023) with $\gamma=5.0$.
-We conduct our training in mixed-precision fp16.
+All training is conducted in mixed-precision FP16.
 Hardware.
-All the codes are implemented in Pytorch utilizing Diffusers (von Platen et al.
-2022) framework and we only conduct training on a single A100 40GB in both two stages.
-The training takes around 3 days for stage 1 on VITON-HD (Choi et al.
-2021) and around 30 hours for stage 2 on both VVT and TikTokDress.
-Additionally, the FPS benchmarks reported in the paper are also obtained using the A100 GPU.
+The code is implemented in PyTorch using the Diffusers (von Platen et al.
+2022) framework, and training is performed on a single A100 GPU with 40GB memory for both stages.
+The training for stage 1 on VITON-HD (Choi et al.
+2021) takes approximately 3 days, while stage 2 on both the VVT and TikTokDress datasets takes around 30 hours.
+The FPS benchmarks reported in the paper were also measured using the A100 GPU.
 Training.
 In the first stage, we optimize the Main UNet decoder, Garment UNet, and Pose Encoder on VITON-HD (Choi et al.
-2021) while other modules are kept fixed.
-The training resolution at this stage remains at $1024\times 768$, consistent with the original resolution.
-Then, we sample a clip of 16 frames with resolution $512\times 384$ for each video as input for stage 2.
-We optimize motion modules only in this stage.
+2021), keeping all other modules fixed.
+The training resolution for this stage is maintained at $1024\times 768$, consistent with the original dataset resolution.
+In the second stage, we sample a clip of 16 frames with a resolution of $512\times 384$ from each video as input.
+At this stage, we optimize only the motion-related modules.
 Appendix B Different Masking Attention Mechanisms
@@
 In the Full Attention configuration ($a$), the attention mask is filled with $0.0$, allowing each frame to attend to all others.
-However, this can lead to less accurate features (i.e., $t+2$, shown in white) influencing the more accurate ones (i.e., $t+1$, shown in red).
-In the Half Attention ($b$) setup, the query frame always attends to the accurate features ($t+1$), enabling the transfer of quality features to less accurate ones.
-The Quarter Attention($c$) strategy allows frames with less accurate features to attend to all other frames while ensuring that the accurate features do not attend to less accurate ones, thereby preserving their quality.
-Finally, Causal Attention ($d$) ensures that the computation for each frame should only depend on the succeeding frames.
-We perform experiments and find out that Half Attention($b$) yields the best performance.
+However, this setup can result in less accurate features (e.g., $t+2$, shown in white) influencing more accurate ones (e.g., $t+1$, shown in red).
+In the Half Attention setup ($b$), the query frame always attends to the accurate features ($t+1$), enabling the transfer of high-quality features to less accurate ones.
+The Quarter Attention setup ($c$) allows frames with less accurate features to attend to all others while ensuring that accurate features do not attend to less accurate ones, thereby preserving their quality.
+Finally, the Causal Attention setup ($d$) ensures that each frame’s computation depends only on succeeding frames (less noise frames).
+Our experiments reveal that Half Attention ($b$) yields the best performance.
 Figure 9: The attention mask used in different masked temporal attention mechanisms in ShiftCaching.
-White cells in the matrix represent a value of 0.0, while gray cells indicate a value of $\infty$- ∞.
+White cells in the matrix represent a value of 00, while gray cells indicate a value of $\infty$- ∞.
 Appendix C TikTokDress Dataset
 Fig.
-12 showcases examples from our dataset, which is diverse in background (ranging from indoor to outdoor settings), camera positions (from top to bottom angles), person distances (from far to close to the camera), skin tones (including black, yellow, and white), and garment types (such as t-shirts, sweaters, crop tops, long-sleeve t-shirts, and dresses).
-Based on this diversity, we believe that our dataset better reflects real-life scenarios for virtual try-ons.
+12 showcases examples from our dataset, highlighting its diversity in backgrounds (ranging from indoor to outdoor settings), camera positions (including top and bottom angles), person distances (ranging from far to close to the camera), skin tones (spanning Black, Asian, and White), and garment types (such as T-shirts, sweaters, crop tops, long-sleeve T-shirts, and dresses).
+This diversity makes our dataset more representative of real-life scenarios for virtual try-on applications.
 Figure 10: Impact of Mask Quality on Try-On Results.
-The first row shows target clothing images.
-The second row demonstrates the effects of an inaccurate mask, leading to flawed try-on results in the third row.
-The fourth row highlights how a precise mask significantly improves quality and accuracy.
-Accurate masks are crucial for high-quality try-ons.
+The first row displays the target clothing images.
+The second row illustrates the effects of an inaccurate mask, resulting in flawed try-on outputs shown in the third row.
+The fourth row demonstrates how a precise mask significantly enhances the quality and accuracy of the results.
+Accurate masks are essential for achieving high-quality try-on outcomes.
 Figure 11: Examples of T-pose
-Agnostic masks are crucial for achieving accurate try-on results.
-For example, in Fig.
-10, an incorrect agnostic mask (highlighted by the red circle) causes the model to hallucinate a garment to fit the mask, resulting in a try-on outcome where the garment appears as a tank top (highlighted by the blue rectangle in the first row).
-With a correct garment mask, a good try-on result is achieved (highlighted by the green rectangle in the first row).
-In the second row, a garment that is not fully masked leads to the model hallucinating additional red stripes in the middle of the red garment (highlighted by the blue rectangle in the second row).
-Again, a correct garment mask produces a good try-on result (highlighted by the green rectangle in the second row).
+Agnostic masks are essential for achieving accurate try-on results.
+For instance, in Fig.
+10, an incorrect agnostic mask (highlighted by the red circle) prompts the model to hallucinate a garment to fit the mask, resulting in a try-on outcome where the garment appears as a tank top (highlighted by the blue rectangle in the first row).
+Conversely, using a correct garment mask produces a satisfactory try-on result (highlighted by the green rectangle in the first row).
+Similarly, in the second row, a garment that is not fully masked causes the model to hallucinate additional red stripes in the middle of the red garment (highlighted by the blue rectangle in the second row).
+Once again, a correct garment mask leads to an accurate try-on result (highlighted by the green rectangle in the second row).
 Figure 12: Examples from our dataset, which feature diverse poses and high-quality garments.
@@
 2024) for both garment and arm segmentation.
-Given a video with multiple frames showing diverse poses, we select the frame that displays as many visible body parts as possible and is clear, similar to Fig.
+From a video with multiple frames showing diverse poses, we select the frame that displays the most visible body parts and has the highest clarity, as illustrated in Fig.
 11.
-We apply the algorithm described in Algorithm 1 for this frame selection.
-The agnostic masks are a combination of the garment mask and the arm mask.
-For garment masks, we choose positive points in the prompting process, specifically the body center, right shoulder, and left shoulder.
-The body center is defined as the midpoint between the neck, right hip, and left hip (see points 1, 8, and 11 in Fig.
+We apply the algorithm described in Algorithm 1 to facilitate this frame selection.
+The agnostic masks are created by combining the garment mask and the arm mask.
+For garment masks, we select positive points during the prompting process, specifically the body center, right shoulder, and left shoulder.
+The body center is defined as the midpoint between the neck, right hip, and left hip (refer to points 1, 8, and 11 in Fig.
 11).
-For the negative points, we select noise, the left eye, and the right eye.
-For arm masks, the positive points we choose are the left elbow, right elbow, left wrist, and right wrist.
-For the negative points, we choose the right knee, nose, and left knee.
-After obtaining the garment and arm masks, we combine them and dilate the result three to seven times to ensure it fully conceals the original garment’s shape.
+Negative points include noise, the left eye, and the right eye.
+For arm masks, the positive points selected are the left elbow, right elbow, left wrist, and right wrist, while negative points include the right knee, nose, and left knee.
+After generating the garment and arm masks, we combine them and apply dilation three to seven times to ensure the mask fully conceals the original garment’s shape.
 Garment Image.
 To extract garment images from videos, we use the Google Lens engine to find the best match at the highest possible resolution.
-In cases where Google Lens does not return a satisfactory result, we manually search on Google using keywords related to the type of clothing, color, and brand, selecting the best match.
-If we are unable to find a suitable garment image, we exclude that video from our dataset.
-After obtaining the garment image, in some cases, we horizontally flip the image to align it with the video.
+If Google Lens does not provide a satisfactory result, we manually search on Google using keywords related to the garment type, color, and brand, selecting the most suitable match.
+If no appropriate garment image can be found, we exclude the video from our dataset.
+After obtaining the garment image, we occasionally horizontally flip it to align with the video.
 Algorithm 1 Selecting the Optimal Frame for SAM2 Prompting
@@
 In the video try-on task, we compare our method on the VVT dataset (Dong et al.
-2019b) against ViViD (Fang et al.
-2024) in Fig.
+2019b) with ViViD (Fang et al.
+2024), as shown in Fig.
 13.
-As demonstrated in video $b.mp4$(see in Supplemental Videos), our method consistently preserves the garment texture throughout the entire video better than ViViD in maintaining texture stability over time.
-In video $d.mp4$ (see in Supplemental Videos), the “GAP” letters on the garment in ViViD exhibit noticeable flickering, a result of its temporal averaging technique.
-In contrast, our method, which employs the ShiftCaching technique, significantly reduces this flickering, ensuring smoother and more stable visual results (Notes: we recommend to see in the supplementary videos).
-Figure 13: Qualitative comparison between our method and ViViD. Our model achieved robust results through long video, thereby forming a coherent video sequence.
-Figure 14: More examples of our video try-on results on our TikTokDress dataset.
-The first row illustrates how well our method’s performs in various backgrounds.
-The second row highlights our method’s ability to preserve texture consistency over time.
+In video $b.mp4$ (see Supplemental Videos), our method consistently preserves garment textures throughout the video, outperforming ViViD in maintaining texture stability over time.
+Similarly, in video $d.mp4$ (see Supplemental Videos), the”GAP” letters on the garment produced by ViViD exhibit noticeable flickering due to its temporal averaging technique.
+In contrast, our method, which incorporates the ShiftCaching technique, significantly reduces flickering, ensuring smoother and more stable visual results.
+We recommend viewing the supplementary videos for a detailed comparison.
+Figure 13: Qualitative comparison between our method and ViViD. Our model achieves robust results over long videos, producing a coherent video sequence.
+Figure 14: Additional examples of our video try-on results on the TikTokDress dataset.
+The first row illustrates how effectively our method performs across various backgrounds, while the second row highlights its ability to preserve texture consistency over time.
 Qualitative results on the TikTokDress dataset.
@@
 14 showcases additional results of our video virtual try-on method on the TikTokDress dataset.
-By utilizing SAM 2 (Ravi et al.
-2024) and manually refining masks on our dataset, combining with the SAM2 frame selection algorithm Algorithm 1, our try-on results demonstrate superior performance compared to those trained on the VVT dataset.
-Our method ensures consistent texture quality for garments and, when combined with ShiftCaching, effectively produces continuous and smooth long videos (Notes: we recommend to see more detail in the supplementary videos).
+By leveraging SAM 2 (Ravi et al.
+2024) and manually refining masks in our dataset, combined with the SAM 2 frame selection algorithm (Algorithm 1), our try-on results demonstrate superior performance compared to those trained on the VVT dataset.
+Our method ensures consistent garment texture quality and, when paired with ShiftCaching, effectively generates continuous and smooth long videos.
+More details can be found in the supplementary videos).